Circular No. 8747

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(134340) PLUTO, (136199) ERIS, AND (136199) ERIS I (DYSNOMIA)

Following the Aug. 24 resolution by the IAU to the effect that the solar system contains eight “planets” (Mercury–Neptune), with (1) Ceres, Pluto (cf. IAUC 255), and 2003 UB_{313} (cf. IAUC 8577) to be considered representative “dwarf planets”, the Minor Planet Center included Pluto and 2003 UB_{313} (along with two other new potential dwarf-planet candidates) in the standard catalogue of numbered objects with well-determined orbits as (134340) and (136199), respectively (see MPC 57525). Following near-unanimous acceptance by both the Committee on Small-Body Nomenclature and the Working Group on Planetary-System Nomenclature (in consultation with the discovery team), the IAU Executive Committee has now approved the names Eris for (136199) and Dysnomia for its satellite (136199) Eris I [formerly S/2005 (2003 UB_{313}) 1; cf. IAUC 8610].

COMET 178P/HUG-BELL Comet P/1999 X1 = 2006 O1 (cf. IAUC 8730) has been given the

permanent number 178P (cf. MPC 57382).

COMET P/2006 R1 (SIDING SPRING) Additional astrometry and the following improved orbital elements for

this comet (cf. IAUC 8744) appear on MPEC 2006-R47. The orbital period, P, is the shortest known for a comet with a retrograde orbit.

T = 2006 Sept. 3.833 TT  ω = 249.^{o}322
e = 0.68901  Ω = 218.561 }2000.0
q = 1.66976 AU  i = 160.021
a = 5.36919 AU n^{o} = 0.079221 P = 12.4 years

SUPERNOVA 2006ev IN UGC 11758 Michel Ory, Vicques, Switzerland, reports the discovery of an appar-

ent supernova (red mag ≈ 16.6) on CCD images taken on Sept. 12.9 and 13.8 UT with a 0.61-m f/3.88 reflector. The new object is located at α = 21^{h}30^{m}59^{s}.26 ± 0^{s}.05, δ = +13^{o}59′21′′.2 ± 0′′.2 (equinox 2000.0), which is 23′′ east and 11′′ north of the nucleus of UGC 11758. Nothing is visible at this location on a red Palomar Sky Survey plate from 1953 Aug. 19 or a red U.K. Schmidt Telescope plate from 1991 Aug. 14.

2006 September 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green